OpenStudy (anonymous):

Is there a chain rule method to differentiate sin(2x), rather than using the conventional 2sinxcosx???

OpenStudy (anonymous):

\[\frac{ dsin(2x) }{ dx }=\frac{ dsin(2x) }{ d(2x) }\frac{ d(2x) }{ dx}=\cos(2x)·2\]

OpenStudy (anonymous):

I think the way CarlosGP did it is usually considered the "conventional" way.

OpenStudy (anonymous):

which is, in fact, the chain rule. Starting with sin(2x)=2sin(x)cos(x) is a bit more complex and unnecessary. It is like finding the derivative of sin(x) through sin(x)=sqrt[(1-cos(2x))/2]

OpenStudy (anonymous):

its really awkward to read!

OpenStudy (anonymous):

how 2 xpand ur ans vertically??

OpenStudy (anonymous):

Oh....got..it!!

OpenStudy (anonymous):

TNX